OpenStudy (anonymous):

(a + b)(a^2 - ab + b^2) i got a^2-ab^3?

OpenStudy (anonymous):

(a + b)(a^2 - ab + b^2) =a^3-a^2*b+B^2*a+a^2*b-a*b^2+b^3 =(a^3+b^3) the middle terms will cancel

OpenStudy (anonymous):

\[(a+b)(a^2-ab+b^2)=a^3-a^2b+ab^2+a^2b-ab^2+b^3= a^3+b^3\]

OpenStudy (anonymous):

ok i have one more: (x^2 + 2x - 1)(x^2 + 2x + 5)